# Persona-driven Simulation of Voting Behavior in the European Parliament with Large Language Models

### v2

@@ 5.2 Default Model Bias
In addition to prompting LLMs to adopt the identity of a specific MEP, we also prompt LLMs directly with the default system prompt "You are a helpful assistant." and reasoning chains. We find that both Llama3-70B and Qwen-72B vote in favor of all but four proposals, while Llama3-8B votes in favor of all but one proposal. This voting behavior aligns most with the group lines of S&D and Renew. Qwen-7B votes in favor of 39 of 47 proposals, abstains in five and votes against in three which aligns closest to Greens/EFA. This behavior is in line with previous research rutinowski2024selfperception; santurkar2023opinions that found a progressive and left-centered bias in many large language models.
@@ 6.1 Counterfactual Debates
One issue that could cause the model to disproportionally predict votes in favor of a proposal may stem from the way the content of the proposals is introduced (cf. section 4.2). When prompting Llama3-70B to determine the stance of each speech that is used to inform about a proposal, the model classifies nearly $67\%$ of speeches as being in favor and only about $32\%$ as being against the respective proposal (roughly $1\%$ are classified as neutral). The fact that the majority of the speeches frame proposals positively could lead to the model being biased toward predicting votes in favor of a proposal.
To control for this effect, we replace the actual debates with counterfactual ones. We instruct Llama3-70B to generate speeches that take the opposite stance of the original speeches, and prompt it to only use information provided in the original speech. To verify whether the generated counterfactual speeches accurately reflect the opposite stance, three voluntary annotators rate a randomly chosen speech for each of the 47 proposals. Based on a majority vote, 43 out of 47 speeches indeed align with the opposite stance relative to the original speeches, with a high level of agreement among annotators (Fleiss’ Kappa score of $0.800$).
When now prompting the models without personas the number of votes in favor of proposals reduces from 43 to 32 for Llama3-70B and from 46 to 38 for Llama3-8B, showing that the debate context can have limited influence on the final vote choice for these models. On the other hand voting behavior for both Qwen-72B and Qwen-7B does not change.
For computational reasons we restrict our experiments with counterfactual speeches with personas to attribute prompting and Llama3-70B with the no reasoning approach. We observe that using counterfactual speeches to describe the proposals decreases the prediction performance to a weighted F1 score of $0.753$. However, this still outperforms all Llama3-8B approaches with the original speeches (cf. Table 1), indicating that the counterfactual debates still provide general information about the proposals.
Figure 3 shows that prompting the model with counterfactual speeches indeed removes the FOR bias and overall even slightly introduces an AGAINST bias. However, this effect is not uniform across groups. While far-left and far-right groups (GUE/NGL, ECR, ID) vote AGAINST significantly more often, the voting behavior of centrist groups (S&D, Renew, Greens/EFA) remains largely unchanged. This suggests that personas continue to play a substantial role in the final vote prediction, despite the stance expressed in the speeches. Another group, that votes AGAINST significantly more often is the group of Non-Inscrits (NI), who are not part of any major political group.
@@ 6.2 Influence of different attributes
As we create personas for politicians of public interest, the models will likely have information about them in their training data. We are interested in if the model can make good predictions even when only giving the name of the politician in context. If additional attributes improve performance, we want to find which ones improve prediction capability the most. Therefore, for each MEP, we prompt the model with (i) only the name, (ii) the name and a single attribute, and (iii) the name and all attributes. To reduce computation cost, we restrict these simulations to Llama3-8B with reasoning, as Qwen-7B showed worse performance with attributes. Prompting with only the name of the MEP (weighted F1 score 0.681) already performs marginally better than the majority baseline (0.666). Prompting with all attributes has the best performance (0.728). Incorporating only the name and the national party (0.722) or the group (0.718) into the persona prompt results in the most significant improvement compared to prompting with the name only and nearly matches the effectiveness of prompting with all attributes.
Figure 4 shows that prompting with all attributes compared to providing only the name significantly steers the model towards a more accurate simulation for groups of the political edge, but also incorrectly steers personas of center left parties to vote in favor of even more proposals.
@@
In addition to analyzing the votes cast by the LLM, we analyze the reasoning chains it provides to justify the voting decision. Specifically, we investigate which persona attributes are explicitly mentioned in the reasoning chains to estimate how much the LLM focuses on the characteristics of an individual persona (cf. Appendix E.1 for details). We emphasize that the reasoning chains should not be interpreted as the LLMs’ internal decision making process, as they may significantly misrepresent the actual reasons behind the predictions of a model turpin2023dontsay. However, we hypothesize that the presence of certain attributes in the reasoning chains may still influence the voting decisions, as transformer models generate content in an autoregressive manner.
As shown in Table 2, the attributes provided in the prompt also appear more often in the reasoning chains. When all attributes are given, the national party is mentioned most often, followed by the European group and the country. Considering this, along with the fact that using only the national party along with the name achieves almost the same performance as the entire set of attributes, this supports the argument of current political research, which indicates that the primary principal of MEPs is their national party willumsen2023policy.
The results in Table 2 for prompting with only the name further indicate that the model has knowledge about MEPs to some extent. For 150 of 653 MEPs in groups, the LLM mentioned the correct European group in the reasoning chains at least once. We assume that the pre-training data of the model contains more information than the information we provide in the prompt. However, prediction results still improve when using all attributes for these politicians, compared to using only the name, indicating that persona attributes have an additional steering effect, even when the pre-training data of the model contains information about the person.
@@
In addition, we evaluate whether the model takes decisions solely based on either the national party or the group. For this we calculate the median variance of both predicted and actual votes per group and national party across all proposals to check if the model’s predictions are always the same for personas of the same national party or European group. The median variance is $0.11$ for national parties and $0.08$ for European groups. These values are only slightly lower than the median variance of actual votes, which are $0.13$ and $0.10$, respectively. The model simulates slightly more party cohesion than in reality, but not full cohesion.
@@ 7 Discussion
We see that all models struggle to predict ABSTETNTION votes correctly. We hypothesize that this could be due to the following three reasons. First, only around $1\%$ of speeches that we leverage as descriptors for the different proposals are classified as neutral by Llama3-70B. We suggest that incorporating more speeches with a neutral stance (i.e. neither advocating FOR or AGAINST a proposal) could increase the number of ABSTENTION votes.
@@
Second, our simulation setup does not take the strategic aspects of coalitions into account. We find that the majority of actual abstention votes are cast by MEPs from groups that form the opposition on most proposals. Groups that form coalitions cannot allow abstentions if they want to retain legislative power, whereas opposition groups have more leeway willumsen2012strategic.
@@
Third, unlike real politicians, our models face no consequence for voting against the group line. Real politicians often have competing interests, which can explain why they may abstain from voting. For example, when the position of the European group conflicts with the national party, politicians are caught in a dilemma. They must balance loyalty to the European group, important for career advancement and influence, with loyalty to the national party, which controls the selection of candidates for future elections muhlbock2017legislators. We assume that LLMs do not capture these complex dynamics and, as a result, are unlikely to take them into account when making predictions.
@@
In addition to determining overall prediction performance of LLMs we find interesting takeaways for further persona based simulations. rottger2024political have shown that models produce different answers on the Politicial Compass Test, when they respond in an open format before being forced to choose a multiple choice option. Conversely, in our approach we have ground-truth data and can see that this open reasoning format improves performance across almost all models tested, especially for personas, that do not have the same voting behavior as the base model. For Qwen-7B it steers the model more, leading to better ABSTENTION and AGAINST performance. We propose that this should be researched further to see if it holds across different contexts.
@@
When exposed to counterfactual debates, LLMs with politically divergent personas (those whose voting behavior differs most from the model’s default system prompt voting behavior, e.g. politicians from ECR or ID) change their predicted voting behavior much more significantly than LLMs with politically aligned personas (those whose voting behavior matches the base model). Similar to bias in multi-agent Debates taubenfeld2024systematic this points towards an inability of LLMs to consistently replicate the viewpoints of politically divergent personas as their fundamental position on a topic can be changed by counterfactual arguments, whereas LLMs prompted with politically aligned personas do not change their fundamental position as easily. However, as our setup has many confounding factors, we call for further research in a more controlled environment to test this hypothesis.
@@ 8 Conclusion
We simulate individual voting behavior of MEPs by prompting LLMs with personas and set a baseline for vote prediction of politicians in the European Parliament. Our results align with current political research, which suggests that the national party is the primary principal for MEPs willumsen2023policy. We find further evidence for a center-left leaning bias of LLMs both in vote prediction and in staying true to the persona’s position, when presented with counterfactual debates. We find that persona-driven LLMs are more steerable when responding in open text format before answering a multiple choice question. This suggests that future work using personas should prioritize open-ended reasoning, particularly when employing smaller models. Additionally, we find that for public figures like politicians, that likely are part of the training data of LLMs, prompting with additional information can help steer the model more. To support further research in political simulation, we release our personas and simulation code. We hope the personas can facilitate further research in other directions and because our approach works with easily obtainable attribute data and general debates, it can work for other political systems as well.
@@
As our computation capacity is limited we had to select a subset of all roll-call votes. Therefore we determined that roll-call votes, that had their own official press release and a debate solely focused on their topic, would be the most relevant and important votes. This, however, can miss certain important votes, e.g. a number of proposals that were taken to reduce and monitor migration and asylum in the EU. None of these roll-call votes had their own press release or a standalone debate. Instead, there was one joint debate for the whole Migration and Asylum package. Therefore, these roll-call votes are not part of our experiments.
@@
As there is a multitude of languages spoken in the European Parliament and the majority of languages are not supported by current LLMs, we use machine translation for the speeches. This can introduce a bias in our debates and inaccurately reflect the viewpoint of certain countries. However, we argue that excluding speeches of low-resource languages would introduce an even greater bias.
@@
As we do not have access to the training data of the models, we can not know if the voting results appear in the pre-training data of the models. We rely on information given publicly about the data cutoff date by the creators, which we can not verify ourselves.
@@
Additionally, while these specific votes are after the pre-training cutoff date, there can be votes about similar proposals, that can be part of the pre-training data.
@@
We are aware that using LLMs to simulate democratic processes raises ethical questions in itself. Our primary goal is to see how well personas can eliminate political bias inherent to LLMs and steer LLMs towards better predictions. We do not propose to use LLMs for actual political decision making. Here, we focus on ethical problems, which we encountered during our project that we find particularly problematic. First, persona prompted LLMs can be used to impersonate real people and spread false information about politicans’ motives to vote a certain way. Especially with reasoning prompts we see arguments, that contain information directly from the politician’s life and sound plausible, but argue for a position the politician did not take in the vote. Second, simply prompting an aligned LLM to create a counterfactual speech arguing against a real political speech can create harmful content, e.g. giving a LLM the task to create a speech of a politician arguing for banning forced labor products can create a speech promoting forced labor as the economically responsible thing to do. Additionally, these impersonated speeches generated by LLMs, based on real politicians, could be used to hijack democratic processes and spread false information. Debates could be reformulated by LLMs into speeches that are expected of a certain party/politician, which formulate different points and steer the discussion away from real arguments.
@@ Appendix A Groups of the European Parliament
The European Parliament in our voting period consists of seven political groups and a set of Non-Inscrits (NI), who are not part of a major political group. The groups, ordered from political left-leaning to political right leaning are: The Left group in the European Parliament (GUE/NGL), Group of the Progressive Alliance of Socialists and Democrats in the European Parliament (S&D), Group of the Greens/European Free Alliance (Greens/EFA), Renew Europe Group (Renew), Group of the European People’s Party (EPP), European Conservatives and Reformists Group (ECR) and Identity and Democracy Group (ID). Note that this left to right ordering does not strictly hold true on all issues. We display exact group lines for each group and vote in our simulation in Table 6.
@@ Appendix B Performance per Group
We show the prediction performance of our simulation for each group when simulating with Llama3-70B and with Qwen-72B in Table 3. Across all methods the votes of center-left and progressive parties are predicted best (S&D, Renew), while votes of groups of the political edge (GUE/NGL, ECR, ID) are predicted worst. We see that personas can influence voting behavior of LLMs, but similar to previous research cannot achieve high prediction performance for European groups with which the base model is not well aligned with chalkidis2024investigating.
Llama3-70B Overall GUE/NGL S&D Greens/EFA Renew EPP ECR ID NI all attributes (r) 0.793 0.711 0.924 0.867 0.900 0.804 0.592 0.529 0.596 wikipedia (r) 0.779 0.669 0.919 0.849 0.894 0.795 0.543 0.538 0.602 all attributes (nr) 0.772 0.671 0.911 0.804 0.899 0.792 0.529 0.560 0.569 wikipedia (nr) 0.770 0.662 0.924 0.854 0.895 0.786 0.507 0.528 0.588 Qwen-72B Overall GUE/NGL S&D Greens/EFA Renew EPP ECR ID NI all attributes (r) 0.789 0.735 0.917 0.842 0.899 0.799 0.574 0.574 0.563 wikipedia (r) 0.773 0.703 0.918 0.855 0.892 0.787 0.502 0.548 0.576 all attributes (nr) 0.773 0.687 0.912 0.846 0.900 0.777 0.516 0.578 0.525 wikipedia (nr) 0.760 0.664 0.915 0.849 0.899 0.771 0.475 0.523 0.559 majority baseline 0.666 0.521 0.909 0.768 0.901 0.762 0.283 0.135 0.413
@@
Additionally, we show the voting behavior of the best-performing Qwen approach with Wikipedia personas and reasoning in Figure 5. Compared to the best Llama3-70B approach, the model does predict ABSTENTIONS more often, but also struggles with predicting them correctly, as both Precision (0.129) and Recall (0.095) are low for ABSTENTION votes.
@@
Appendix C Counterfactual Debates

### v3

@@ 5.2 Default Model Bias
In addition to prompting LLMs to adopt the identity of a specific MEP, we also prompt LLMs directly with the default system prompt "You are a helpful assistant." and reasoning chains. We find that both Llama3-70B and Qwen-72B vote in favor of all but four proposals, while Llama3-8B votes in favor of all but one proposal. This voting behavior aligns most with the group lines of S&D and Renew. Qwen-7B votes in favor of 39 of 47 proposals, abstains in five, and votes against in three, which aligns closest to Greens/EFA. This behavior is in line with previous research Rutinowski et al. (2024); Santurkar et al. (2023) that found a progressive and left-centered bias in many large language models.
@@ 6.1 Counterfactual Debates
One issue that could cause the model to disproportionately predict votes in favor of a proposal may stem from the way the content of the proposals is introduced (cf. section 4.2). When prompting Llama3-70B to determine the stance of each speech used to inform about a proposal, the model classifies nearly $67\%$ of speeches as being in favor and only about $32\%$ as being against the respective proposal (roughly $1\%$ are classified as neutral). The fact that the majority of the speeches frame proposals positively could lead to the model being biased toward predicting votes in favor of a proposal.
To control for this effect, we replace the actual debates with counterfactual ones. We instruct Llama3-70B to generate speeches that take the opposite stance of the original speeches and prompt it to use only the information provided in the original speech. To verify whether the generated counterfactual speeches accurately reflect the opposite stance, three voluntary annotators rate a randomly chosen speech for each of the 47 proposals. Based on a majority vote, 43 out of 47 speeches indeed align with the opposite stance relative to the original speeches, with a high level of agreement among annotators (Fleiss’ Kappa score of $0.800$).
When prompting the models without personas, the number of votes in favor of proposals reduces from 43 to 32 for Llama3-70B and from 46 to 38 for Llama3-8B, showing that the debate context can have a limited influence on the final vote choice for these models. On the other hand, voting behavior for both Qwen-72B and Qwen-7B does not change.
For computational reasons, we restrict our experiments with counterfactual speeches with personas to attribute prompting and Llama3-70B using the no reasoning approach. We observe that using counterfactual speeches to describe the proposals decreases the prediction performance to a weighted F1 score of $0.753$. However, this still outperforms all Llama3-8B approaches with the original speeches (cf. Table 1), indicating that the counterfactual debates still provide general information about the proposals.
Figure 3 shows that prompting the model with counterfactual speeches indeed removes the FOR bias and overall slightly introduces an AGAINST bias. However, this effect is not uniform across groups. While far-left and far-right groups (GUE/NGL, ECR, ID) vote AGAINST significantly more often, the voting behavior of centrist groups (S&D, Renew, Greens/EFA) remains largely unchanged. This suggests that personas continue to play a substantial role in the final vote prediction, despite the stance expressed in the speeches. Another group that votes AGAINST significantly more often is the group of Non-Inscrits (NI), who are not part of any major political group.
@@ 6.2 Influence of different attributes
As we create personas for politicians of public interest, the models will likely have information about them in their training data. We are interested in whether the model can make good predictions even when only given the name of the politician in context. If additional attributes improve performance, we want to determine which ones enhance prediction capability the most. Therefore, for each MEP, we prompt the model with (i) only the name, (ii) the name and a single attribute, and (iii) the name and all attributes. To reduce computation costs, we restrict these simulations to Llama3-8B with reasoning, as Qwen-7B showed worse performance with attributes. Prompting with only the name of the MEP (weighted F1 score 0.681) already performs marginally better than the majority baseline (0.666). Prompting with all attributes yields the best performance (0.728). Incorporating only the name and the national party (0.722) or the group (0.718) into the persona prompt results in the most significant improvement compared to prompting with the name only, and it nearly matches the effectiveness of prompting with all attributes.
Figure 4 shows that prompting with all attributes, compared to providing only the name, significantly steers the model towards a more accurate simulation for groups on the political edge, but also incorrectly steers personas of center left parties to vote in favor of even more proposals.
@@
In addition to analyzing the votes cast by the LLM, we analyze the reasoning chains it provides to justify the voting decision. Specifically, we investigate which persona attributes are explicitly mentioned in the reasoning chains to estimate how much the LLM focuses on the characteristics of an individual persona (cf. Appendix G.1 for details). We emphasize that the reasoning chains should not be interpreted as the LLMs’ internal decision making process, as they may significantly misrepresent the actual reasons behind the predictions of a model Turpin et al. (2023). However, we hypothesize that the presence of certain attributes in the reasoning chains may still influence the voting decisions, as transformer models generate content in an autoregressive manner.
As shown in Table 2, the attributes provided in the prompt also appear more often in the reasoning chains. When all attributes are given, the national party is mentioned most often, followed by the European group and the country. Considering this, along with the fact that using only the national party along with the name achieves almost the same performance as the entire set of attributes, this supports the argument of current political research, which indicates that the primary principal of MEPs is their national party Willumsen (2023).
The results in Table 2 for prompting with only the name further indicate that the model has knowledge about MEPs to some extent. For 150 of 653 MEPs in groups, the LLM mentioned the correct European group in the reasoning chains at least once. We assume that the model’s pre-training data contains more information than the information we provide in the prompt. However, prediction results still improve when using all attributes for these politicians compared to using only the name, indicating that persona attributes have an additional steering effect, even when the model’s pre-training data contains information about the person.
@@
In addition, we evaluate whether the model takes decisions solely based on either the national party or the group. For this, we calculate the median variance of both predicted and actual votes per group and national party across all proposals to check if the model’s predictions are always the same for personas of the same national party or European group. The median variance is $0.11$ for national parties and $0.08$ for European groups. These values are only slightly lower than the median variance of actual votes, which are $0.13$ and $0.10$, respectively. The model simulates slightly more party cohesion than in reality, but not full cohesion.
@@ 7 Discussion
We see that all models struggle to predict ABSTENTION votes correctly. We hypothesize that this could be due to the following three reasons. First, only around $1\%$ of speeches that we leverage as descriptors for the different proposals are classified as neutral by Llama3-70B. We suggest that incorporating more speeches with a neutral stance (i.e., neither advocating FOR nor AGAINST a proposal) could increase the number of ABSTENTION votes.
@@
Second, our simulation setup does not take the strategic aspects of coalitions into account. We find that the majority of actual abstention votes are cast by MEPs from groups that form the opposition on most proposals. Groups that form coalitions cannot allow abstentions if they want to retain legislative power, whereas opposition groups have more leeway Willumsen and Öhberg (2012).
@@
Third, unlike real politicians, our models face no consequences for voting against the group line. Real politicians often have competing interests, which can explain why they may abstain from voting. For example, when the position of the European group conflicts with the national party, politicians are caught in a dilemma. They must balance loyalty to the European group, important for career advancement and influence, with loyalty to the national party, which controls the selection of candidates for future elections Mühlböck and Yordanova (2017). We assume that LLMs do not capture these complex dynamics and, as a result, are unlikely to take them into account when making predictions.
@@
In addition to determining the overall predictive performance of LLMs, we find interesting takeaways for further persona based simulations. Röttger et al. (2024) have shown that models produce different answers on the Politicial Compass Test, when they respond in an open format before being forced to choose a multiple choice option. Conversely, in our approach, we have ground-truth data and can see that this open reasoning format improves performance across almost all models tested, especially for personas that do not have the same voting behavior as the base model. For Qwen-7B, it steers the model more, leading to better ABSTENTION and AGAINST performance. We argue that this should be kept in mind for further experiments.
@@
When exposed to counterfactual debates, LLMs with politically divergent personas (those whose voting behavior differs most from the model’s default system prompt voting behavior, e.g., politicians from ECR or ID) change their predicted voting behavior much more significantly than LLMs with politically aligned personas (those whose voting behavior matches the base model). Similar to findings in previous work Taubenfeld et al. (2024); Liu et al. (2024), this points towards an inability of LLMs to consistently replicate the viewpoints of politically divergent personas, as their fundamental position on a topic can be changed by counterfactual arguments, whereas LLMs prompted with politically aligned personas do not change their fundamental position as easily. However, stylistic differences beyond stance, such as speech length, can be confounders in our setup, and we call for further experiments in a controlled setup.
@@ 8 Conclusion
We simulate the individual voting behavior of MEPs by prompting LLMs with personas and set a baseline for vote prediction of politicians in the European Parliament. Our results align with current political research, which suggests that the national party is the primary principal for MEPs Willumsen (2023). We find further evidence of a center-left leaning bias of LLMs, both in vote prediction and in staying true to the persona’s position when presented with counterfactual debates. We find that persona-driven LLMs are more steerable when responding in open text format before answering a multiple choice question. This suggests that future work using personas should prioritize open-ended reasoning, particularly when employing smaller models. Additionally, we find that for public figures like politicians, who are likely part of the training data of LLMs, prompting with additional information can help steer the model more effectively. To support further research in political simulation, we release our personas and simulation code. We hope the personas can facilitate further research in other directions, and because our approach works with easily obtainable attribute data and general debates, it can be effective for other political systems as well.
@@
As our computational capacity is limited, we had to select a subset of all roll-call votes. Therefore, we determined that roll-call votes that had their own official press release and a debate solely focused on their topic would be the most relevant and important votes. This, however, may miss certain important votes, e.g., a number of proposals that were introduced to reduce and monitor migration and asylum in the EU. None of these roll-call votes had their own press release or a standalone debate. Instead, there was one joint debate for the entire Migration and Asylum package. Therefore, these roll-call votes are not part of our experiments.
@@
As there is a multitude of languages spoken in the European Parliament and the majority of these languages are not supported by current LLMs, we use machine translation via google translate for the speeches. This can introduce bias into our debates and inaccurately reflect the viewpoints of certain countries. However, we argue that excluding speeches in low-resource languages would introduce an even greater bias.
@@
As we do not have access to the training data of the models, we cannot know if the voting results appear in the pre-training data of the models. We rely on information publicly provided about the data cutoff date by the creators, which we cannot verify ourselves.
@@
Additionally, while these specific votes are after the pre-training cutoff date, there can be votes about similar proposals that can be part of the pre-training data.
@@
We are aware that using LLMs to simulate democratic processes raises ethical questions in itself. Our primary goal is to see how well personas can eliminate the political bias inherent in LLMs and steer LLMs towards better predictions. We do not propose to use LLMs for actual political decision making. Here, we focus on the ethical problems that we encountered during our project, which we find particularly problematic. First, persona prompted LLMs can be used to impersonate real people and spread false information about politicians’ motives to vote a certain way. Especially with reasoning prompts, we see arguments that contain information directly from the politician’s life and sound plausible, but argue for a position the politician did not take in the vote. Second, simply prompting an aligned LLM to create a counterfactual speech arguing against a real political speech can create harmful content; for example, giving an LLM the task of creating a speech for a politician arguing for banning forced labor products can result in a speech promoting forced labor as the economically responsible option. Additionally, these impersonated speeches generated by LLMs, based on real politicians, could be used to hijack democratic processes and spread false information. Debates could be reformulated by LLMs into speeches that are expected from a certain party or politician, which present different points and steer the discussion away from real arguments.
@@ Appendix A Groups of the European Parliament
The European Parliament during our voting period consists of seven political groups and a set of Non-Inscrits (NI), who are not part of a major political group. The groups, ordered from political left-leaning to political right leaning, are: The Left group in the European Parliament (GUE/NGL), the Group of the Progressive Alliance of Socialists and Democrats in the European Parliament (S&D), the Group of the Greens/European Free Alliance (Greens/EFA), the Renew Europe Group (Renew), the Group of the European People’s Party (EPP), the European Conservatives and Reformists Group (ECR), and the Identity and Democracy Group (ID). Note that this left to right ordering does not strictly hold true for all issues. We display exact group lines for each group and vote in our simulation in Table 5.
@@ Appendix B Performance per Group
We show the prediction performance of our simulation for each group when simulating with Llama3-70B and Qwen-72B in Table 3. Across all methods, the votes of center-left and progressive parties are predicted best (S&D, Renew), while the votes of groups on the political edge (GUE/NGL, ECR, ID) are predicted worst. We see that personas can influence the voting behavior of LLMs, but similar to previous research, they cannot achieve high prediction performance for European groups with which the base model is not well aligned Chalkidis (2024).
Llama3-70B Overall GUE/NGL S&D Greens/EFA Renew EPP ECR ID NI all attributes (r) 0.793 0.711 0.924 0.867 0.900 0.804 0.592 0.529 0.596 wikipedia (r) 0.779 0.669 0.919 0.849 0.894 0.795 0.543 0.538 0.602 all attributes (nr) 0.772 0.671 0.911 0.804 0.899 0.792 0.529 0.560 0.569 wikipedia (nr) 0.770 0.662 0.924 0.854 0.895 0.786 0.507 0.528 0.588 Qwen-72B Overall GUE/NGL S&D Greens/EFA Renew EPP ECR ID NI all attributes (r) 0.789 0.735 0.917 0.842 0.899 0.799 0.574 0.574 0.563 wikipedia (r) 0.773 0.703 0.918 0.855 0.892 0.787 0.502 0.548 0.576 all attributes (nr) 0.773 0.687 0.912 0.846 0.900 0.777 0.516 0.578 0.525 wikipedia (nr) 0.760 0.664 0.915 0.849 0.899 0.771 0.475 0.523 0.559 bert baseline 0.762 0.615 0.903 0.860 0.892 0.823 0.536 0.454 0.413 zero-shot baseline 0.755 0.625 0.885 0.820 0.873 0.772 0.535 0.568 0.413 majority baseline 0.666 0.521 0.909 0.768 0.901 0.762 0.283 0.135 0.413
@@
Furthermore, we show the voting behavior of the best-performing Qwen approach with Wikipedia personas and reasoning in Figure 5. Compared to the best Llama3-70B approach, the model predicts ABSTENTION more often but also struggles to predict them correctly, as both Precision (0.129) and Recall (0.095) are low for ABSTENTION votes.
@@
Appendix C Stronger baselines
To provide a comparison for our method, we established additional baselines using a simplified task. In this setup, we analyze the sentiment of the main spokesperson’s speech for each group. If the spokesperson’s speech is positive, we assign a FOR stance to their entire group. This baseline has an advantage because it relies on the spokesperson’s group identity, which our main approach does not utilize. Since the NI group does not have a spokesperson, we simply predict FOR for the politicians of that group.
We use two approaches for this sentiment analysis: 1. Fine-tuned BERT: We fine-tune bert-base-uncased Devlin et al. (2019) on 15,389 speeches from the years 2021 to 2023. To avoid data leakage, this training data is separate from our prediction dataset (2024). We use the MEP’s voting decision as the ground truth label; for example, if an MEP voted FOR, we label their speech as positive. We train the model for 3 epochs with a batch size of 16 and a learning rate of 2e52e-5. 2. Zero-shot LLM: We use the zero-shot sentiment predictions generated by Llama3-70B, following the same setup described in our counterfactual debate ablation.
The results are shown in the lower part of Table 3. The fine-tuned BERT model achieves a mean F1-score of 0.762, and the zero-shot approach achieves 0.755. Both scores are lower than our best performing method (0.793), even though these baselines rely on the strong signal from the group spokesperson.
Appendix D Statistical Analysis
To test whether the differences in prediction performance in Table 1 are statistically significant, we conduct an aligned rank transform (ART) ANOVA to examine the effects of reasoning, persona prompt (i.e., attribute and Wikipedia prompting), and their interaction on performance (i.e., F1-score) across the four models. For the Qwen models, no effects were significant. In contrast, both Llama models showed a significant interaction between reasoning and prompt type (p < 0.05), while neither factor had a significant main effect. This indicates that the benefit of reasoning differs between attribute and Wikipedia prompts. Specifically, reasoning provides a larger performance improvement for attribute prompting.
Appendix E Counterfactual Debates